OpenStudy (anonymous):

let f(x)=-5x-4 and g(x)=6x-7. find f(x)+g(x)

OpenStudy (michele_laino):

hint: \[\Large f\left( x \right) + g\left( x \right) = \left( { - 5x - 4} \right) + \left( {6x - 7} \right) = ...\]

OpenStudy (anonymous):

x-11?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thanks for your help! :)

OpenStudy (michele_laino):

:)